# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 3.4 Horizontal infinitesimal deformations
Suppose that $X$ is compact. Let $(E,\theta)$ be a Higgs bundle of degree 00 on $X$ which is generically regular semisimple with smooth spectral curve.
@@
  • An infinitesimal deformation $(E^{[1]},\theta^{[1]})$ of $(E,\theta)$ is called horizontal if there exists a relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $S\times X$ such that (i) $(\mathcal{E},\theta_{\mathcal{E}})$ is horizontal, (ii) there exists $s_{0}\in S$ and $v\in T_{s_{0}}S$ such that $\iota_{v}^{\ast}(\mathcal{E},\theta_{\mathcal{E}})\simeq(E^{[1]},\theta^{[1]})$. (See §2.3.2 for the notation.)
@@
Let $\mathcal{M}_{H}$ denote the moduli space of stable Higgs bundles of degree 00 and rank $n$ on $X$. There exists the Hitchin fibration $\Phi_{H}:\mathcal{M}_{H}\to{\boldsymbol{A}}_{X,n}$. We obtain the point $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ corresponding to $(E,\theta)$. Let $\mathcal{U}$ be any simply connected open neighbourhood of $\Phi_{H}(E,\theta)$. According to Lemma 3.24, there exists a morphism of complex manifolds $\Psi:\mathcal{U}\to\mathcal{M}_{H}$ such that (i) $\Phi_{H}\circ\Psi=\mathop{\rm id}\nolimits_{\mathcal{U}}$, (ii) the family of Higgs bundles $\Psi(x)$ $(x\in\mathcal{U})$ is horizontal, (iii) $\Psi(\Phi_{H}(E,\theta))=(E,\theta)$.
@@
Then, we define a subcomplex $\mathfrak{A}^{\bullet}$ of $\mathop{\rm Tot}\nolimits\check{C}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ as follows:
@@
Proof   Let us prove that $\Upsilon_{1}\circ\Upsilon_{0}$ is injective. Let $\Upsilon_{1}\circ\Upsilon_{0}(a)=0$ for $a\in H^{1}(\mathfrak{A}^{\bullet})$. Let $(\varphi_{0},\varphi_{P},\psi_{P})\in\mathfrak{A}^{1}$ be a representative of $a$. Because $\Upsilon_{1}\circ\Upsilon_{0}(a)=0$, we obtain that $\varphi_{0}=0$ and $[\varphi_{P}]=0$ in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta_{P})$ $(P\in D(\theta))$. There exist $\nu_{P}\in H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P}))$ such that $\mathop{\rm ad}\nolimits(\theta_{P})\nu_{P}=\varphi_{P}$. Because we have $\varphi_{P|X_{P}\cap X_{0}}=\mathop{\rm ad}\nolimits\theta_{P}(\psi_{P})$, we obtain $\nu_{P|X_{P}\cap X_{0}}=\psi_{P}$. Hence, we obtain $a=0$ in $H^{1}(\mathfrak{A}^{\bullet})$.
@@
The cohomology class $[a]\in H^{1}(\mathfrak{A}^{\bullet})$ is $(\Upsilon_{1}\circ\Upsilon_{0})^{-1}(\nu)$. Its image in $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is $\iota^{{\tt h}}(\nu)$.
@@ 3.4.2 $C^{\infty}$-description of horizontal infinitesimal deformations
Then, $[\tau]=\iota^{{\tt h}}(\nu)$ holds in $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. In particular, $[\tau]\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$. Conversely, any element of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ is described in this way.
@@ 3.4.3 Lagrangian property of the horizontal subspaces
For any Higgs bundle $(E,\theta)$ of rank $n$ of degree 00 on $X$, let $\omega_{\mathcal{M}_{H},(E,\theta)}$ denote the alternative bilinear form on $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ defined as follows [14, (6.5)]:
@@
Proof   For any $\iota^{{\tt h}}(\nu_{i})\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ $(i=1,2)$, there exists a representative $\tau_{i}\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfying the conditions in Corollary 3.32. On $X\setminus\bigcup X_{P}^{\prime}$, we have $\tau_{i}=\tau_{i}^{1,0}$, and hence $\tau_{1}\wedge\tau_{2}=0$. On $X_{P}$, $\tau_{i}^{1,0}$ (resp. $\tau_{i}^{0,1}$) are symmetric (anti-symmetric) with respect to $C_{P}$. We recall that for a symmetric $(n\times n)$-matrix $A$ and an anti-symmetric $(n\times n)$-matrix $B$, we obtain $\mathop{\rm tr}\nolimits(AB)=0$ because
@@ 4.1.1 Vertical infinitesimal deformations
Proof   Because $\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}}=\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))$, we obtain the first claim. Let $U\subset{\mathbb{C}}$ be a neighbourhood of 00. For any $a\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$, there exists a holomorphic line bundle $\mathcal{L}$ on $U\times\Sigma_{\theta}$ which induces $a$ as the infinitesimal part. Then, the second claim is clear.
Any $a\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ is represented by a $(0,1)$-form $\nu$ such that the support of $\nu$ is compact in $\Sigma_{\theta}\setminus\Sigma_{\theta|D(\theta)}$. We obtain the $C^{\infty}$-section $F^{\prime}_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X^{\circ}$ as in §2.2.7 whose support is compact. It naturally extends to a $C^{\infty}$-section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X$. Because $[\theta,F_{\nu}]=0$, it is a $1$-cocycle in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$. The image $\iota^{{\tt v}}(a)\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is represented by $F_{\nu}$.
@@ 4.1.2 The induced perfect pairings
There exists the following decomposition.
@@
Because $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$ is Lagrangian, we obtain the following induced perfect pairings.
@@ 4.1.4 The semi-flat metric
Note that $\Phi_{H}:\mathcal{M}_{H}^{\prime}\to{\boldsymbol{A}}_{X,n}^{\prime}$ is an algebraic integrable system. For each $s\in{\boldsymbol{A}}_{X,n}^{\prime}$, we obtain the spectral curve $\Sigma_{s}\subset T^{\ast}X$, and $\Phi_{H}^{-1}(s)$ is identified with the moduli space $\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s})$ of holomorphic line bundles on $\Sigma_{s}$ of degree $d=n(n-1)(g_{X}-1)$. By choosing $K_{X}^{1/2}$, we obtain the Hitchin section $\Psi$ of $\Phi_{H}$. By using the Hitchin section, we may regard $\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s})$ as the quotient of $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$ by a lattice. The space $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$ is equipped with the real symplectic form $\omega_{s}$ which induces a positive symmetric bilinear form on $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$. The cohomology class $[\omega_{s}]\in H^{2}(\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s}),{\mathbb{R}})$ is a positive polarization in the sense of [8, §3]. According to [8, Theorem 3.8], we obtain a hyperkähler metric $g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}$, which is called the semi-flat metric. Let $g_{\mathop{\rm sf}\nolimits}$ denote the associated Kähler metric with respect to the natural complex structure of $\mathcal{M}^{\prime}_{H}$ as an open subset of the moduli space of Higgs bundles.
@@
Proof   Let $\omega_{1}$ and $\omega_{2}$ be the real symplectic forms on $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ and $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ as in §2.7. Let us recall the construction of the semi-flat metric at $T_{(E,\theta)}\mathcal{M}$ by following [8, §2 and §3].
@@
We obtain the positive symplectic form $\omega_{\Phi_{H}(E,\theta)}=-\omega_{1}^{\lor}$ on $T_{\Phi_{H}(E,\theta)}{\boldsymbol{A}}_{X,n}$. In this way, we obtain the Kähler form $\omega_{{\boldsymbol{A}}^{\prime}_{X,n}}=\{\omega_{s}\}_{s\in{\boldsymbol{A}}^{\prime}_{X,n}}$ on ${\boldsymbol{A}}^{\prime}_{X,n}$. The decomposition $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))=H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\oplus H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ induces the flat connection on the tangent bundle of $\mathcal{M}_{H}^{\prime}$. They give the special Kähler structure on ${\boldsymbol{A}}^{\prime}_{X,n}$. (See [8, §3].) We obtain the special Kähler metric $g_{3}$ of ${\boldsymbol{A}}^{\prime}_{X,n}$ induced by $\omega_{{\boldsymbol{A}}^{\prime}_{X,n}}$. Let $g_{3,{\mathbb{R}}}$ denote the real part of $g_{3}$.
@@ 4.2.3 Appendix
¯$\overline{\partial}_{E}\tau^{1,0}+[\theta,\tau^{0,1}]=0.$ (38)
@@ 4.3 Harmonic representatives of horizontal infinitesimal deformations
Here, $c_{1}$ is induced by $\Sigma_{\theta}\simeq\Sigma_{t\theta}$, and $c_{2}$ is the morphism $\iota^{{\tt h}}$ in (3.4.1) for $(E,t\theta)$. Let $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. Let us study the harmonic $1$-form ${\tt H}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that $[{\tt H}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$. Let $\|\nu\|$ denote the $L^{2}$-norm of $\nu$ on $\Sigma_{\theta}$. Let $g_{X}$ denote a conformal metric of $X$.
@@ 4.3.1 Asymptotically harmonic representative on $X^{\circ}$
$\left|\sqrt{-1}\int_{N(D)}h_{t}(F_{\nu},F_{\nu})-\sqrt{-1}\int_{\Sigma_{\theta|N(D)}}\nu\wedge\overline{\nu}\right|\\ \leq B_{1}\exp(-B_{2}t)\|\nu\|^{2}.$ (40)
@@
$\sqrt{-1}\int_{\Sigma_{\theta|N(D)}}\nu\wedge\overline{\nu}=\sqrt{-1}\int_{N(D)}\mathop{\rm tr}\nolimits\bigl{(}F_{\nu}\cdot F_{\overline{\nu}}\bigr{)}.$
@@
$\sqrt{-1}\int_{N(D)}h_{t}(F_{\nu},F_{\nu})=\sqrt{-1}\int_{N(D)}\mathop{\rm tr}\nolimits\bigl{(}F_{\nu}\cdot(F_{\nu})^{\dagger}_{h_{t}}\bigr{)}.$
@@
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{g_{X}}((\mathop{\rm tr}\nolimits F_{\nu})\mathop{\rm id}\nolimits_{E})=0$.

### v3

@@ 3.4 Horizontal infinitesimal deformations
Suppose that $X$ is compact. Let $(E,\theta)$ be a Higgs bundle of degree 0 on $X$ which is generically regular semisimple with smooth spectral curve.
@@
  • An infinitesimal deformation $(E^{[1]},\theta^{[1]})$ of $(E,\theta)$ is called horizontal if there exists a relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $S\times X$ such that (i) $(\mathcal{E},\theta_{\mathcal{E}})$ is horizontal, (ii) there exist $s_{0}\in S$ and $v\in T_{s_{0}}S$ such that $\iota_{v}^{\ast}(\mathcal{E},\theta_{\mathcal{E}})\simeq(E^{[1]},\theta^{[1]})$. (See §2.3.2 for the notation.)
@@
Let $\mathcal{M}_{H}$ denote the moduli space of stable Higgs bundles of degree 0 and rank $n$ on $X$. There exist the Hitchin fibration $\Phi_{H}:\mathcal{M}_{H}\to{\boldsymbol{A}}_{X,n}$ and its restriction $\Phi_{H}:\mathcal{M}^{\prime}_{H}\to{\boldsymbol{A}}^{\prime}_{X,n}$. We obtain the point $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ corresponding to $(E,\theta)$. Let $\mathcal{U}$ be any simply connected open neighbourhood of $\Phi_{H}(E,\theta)$ in ${\boldsymbol{A}}^{\prime}_{X,n}$. According to Lemma 3.24, there exists a morphism of complex manifolds $\Psi:\mathcal{U}\to\mathcal{M}^{\prime}_{H}$ such that (i) $\Phi_{H}\circ\Psi=\mathop{\rm id}\nolimits_{\mathcal{U}}$, (ii) the family of Higgs bundles $\Psi(x)$ $(x\in\mathcal{U})$ is horizontal, (iii) $\Psi(\Phi_{H}(E,\theta))=(E,\theta)$.
@@
Then, we define a subcomplex $\mathfrak{A}^{\bullet}$ of $\mathop{\rm Tot}\nolimits\check{C}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ as follows (see Remark 2.17 for $\mathop{\rm Tot}\nolimits\check{C}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$):
@@
Proof  Let us prove that $\Upsilon_{1}\circ\Upsilon_{0}$ is injective. Let $\Upsilon_{1}\circ\Upsilon_{0}(a)=0$ for $a\in H^{1}(\mathfrak{A}^{\bullet})$. Let $(\varphi_{0},\varphi_{P},\psi_{P})\in\mathfrak{A}^{1}$ be a representative of $a$. Because $\Upsilon_{1}\circ\Upsilon_{0}(a)=0$, we obtain that $\varphi_{0}=0$ and $\varphi_{P}]=0$ in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta_{P})$ $(P\in D(\theta))$. There exists $\nu_{P}\in H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P}))$ such that $\mathop{\rm ad}\nolimits(\theta_{P})\nu_{P}=\varphi_{P}$. Because we have $\varphi_{P|X_{P}\cap X_{0}}=\mathop{\rm ad}\nolimits\theta_{P}(\psi_{P})$, we obtain $\nu_{P|X_{P}\cap X_{0}}=\psi_{P}$. Hence, we obtain $a=0$ in $H^{1}(\mathfrak{A}^{\bullet})$.
@@
The cohomology class $a]\in H^{1}(\mathfrak{A}^{\bullet})$ is $(\Upsilon_{1}\circ\Upsilon_{0})^{-1}(\nu)$. Its image in $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is $\iota^{{\tt h}}(\nu)$.
@@ 3.4.2 $C^{\infty}$-description of horizontal infinitesimal deformations
Then, $\tau]=\iota^{{\tt h}}(\nu)$ holds in $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. In particular, $\tau]\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$. Conversely, any element of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ is described in this way.
@@ 3.4.3 Lagrangian property of the horizontal subspaces
For any Higgs bundle $(E,\theta)$ of rank $n$ of degree 0 on $X$, let $\omega_{\mathcal{M}_{H},(E,\theta)}$ denote the alternative bilinear form on $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ defined as follows [14, (6.5)]:
@@
Proof  For any $\iota^{{\tt h}}(\nu_{i})\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ $(i=1,2)$, there exist representatives $\tau_{i}\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfying the conditions in Corollary 3.32. On $X\setminus\bigcup X_{P}^{\prime}$, we have $\tau_{i}=\tau_{i}^{1,0}$, and hence $\tau_{1}\wedge\tau_{2}=0$. On $X_{P}$, $\tau_{i}^{1,0}$ (resp. $\tau_{i}^{0,1}$) are symmetric (anti-symmetric) with respect to $C_{P}$. We recall that for a symmetric $(n\times n)$-matrix $A$ and an anti-symmetric $(n\times n)$-matrix $B$, we obtain $\mathop{\rm tr}\nolimits(AB)=0$ because
@@ 4.1.1 Vertical infinitesimal deformations
Proof  Because $\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}}=\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))$, we obtain the first claim. Let $U\subset{\mathbb{C}}$ be a neighbourhood of 0. For any $a\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$, there exists a holomorphic line bundle $\mathcal{L}$ on $U\times\Sigma_{\theta}$ which induces $a$ as the infinitesimal part. Then, the second claim is clear.
Any $a\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ is represented by a $(0,1)$-form $\nu$ such that the support of $\nu$ is compact in $\Sigma_{\theta}\setminus\Sigma_{\theta|D(\theta)}$. We obtain the $C^{\infty}$-section $F^{\prime}_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X^{\circ}$ as in §2.2.7 whose support is compact. It naturally extends to a $C^{\infty}$-section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X$. Because $\theta,F_{\nu}]=0$, it is a $1$-cocycle in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$. The image $\iota^{{\tt v}}(a)\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is represented by $F_{\nu}$.
@@ 4.1.2 The induced perfect pairings
There exists the following decomposition:
@@
Because $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$ is Lagrangian, we obtain the following perfect pairings induced by $\omega_{\mathcal{M}_{H},(E,\theta)}$.
@@ 4.1.4 The semi-flat metric
Note that $\Phi_{H}:\mathcal{M}_{H}^{\prime}\to{\boldsymbol{A}}_{X,n}^{\prime}$ is an algebraic integrable system. For each $s\in{\boldsymbol{A}}_{X,n}^{\prime}$, we obtain the spectral curve $\Sigma_{s}\subset T^{\ast}X$, and $\Phi_{H}^{-1}(s)$ is identified with the moduli space $\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s})$ of holomorphic line bundles on $\Sigma_{s}$ of degree $d=n(n-1)(g_{X}-1)$. By choosing $K_{X}^{1/2}$, we obtain the Hitchin section $\Psi$ of $\Phi_{H}$. By using the Hitchin section, we may regard $\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s})$ as the quotient of $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$ by a lattice. The space $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$ is equipped with the real symplectic form $\omega_{s}$ which induces a positive symmetric bilinear form on $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$. The cohomology class $\omega_{s}]\in H^{2}(\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s}),{\mathbb{R}})$ is a positive polarization in the sense of [8, §3]. According to [8, Theorem 3.8], we obtain a hyperkähler metric $g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}$, which is called the semi-flat metric. Let $g_{\mathop{\rm sf}\nolimits}$ denote the associated Kähler metric with respect to the natural complex structure of $\mathcal{M}^{\prime}_{H}$ as an open subset of the moduli space of Higgs bundles.
@@
Proof  Let $\omega_{1}$ and $\omega_{2}$ be the real symplectic forms on $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ and $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ as in §2.7. Let us recall the construction of the semi-flat metric at $T_{(E,\theta)}\mathcal{M}_{H}$ by following [8, §2 and §3].
@@
We obtain the positive symplectic form $\omega_{\Phi_{H}(E,\theta)}=-\omega_{1}^{\lor}$ on $T_{\Phi_{H}(E,\theta)}{\boldsymbol{A}}_{X,n}$. In this way, we obtain the Kähler form $\omega_{{\boldsymbol{A}}^{\prime}_{X,n}}=\{\omega_{s}\}_{s\in{\boldsymbol{A}}^{\prime}_{X,n}}$ on ${\boldsymbol{A}}^{\prime}_{X,n}$. The decomposition $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))=H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\oplus H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ induces the flat connection on the tangent bundle of ${\boldsymbol{A}}^{\prime}_{X,n}$. They give the special Kähler structure on ${\boldsymbol{A}}^{\prime}_{X,n}$. (See [8, §3].) We obtain the special Kähler metric $g_{3}$ of ${\boldsymbol{A}}^{\prime}_{X,n}$ induced by $\omega_{{\boldsymbol{A}}^{\prime}_{X,n}}$. Let $g_{3,{\mathbb{R}}}$ denote the real part of $g_{3}$.
@@ 4.2.3 Appendix
$\overline{\partial}_{E}\tau^{1,0}+[\theta,\tau^{0,1}]=0.$ (42)
@@ 4.3 Harmonic representatives of horizontal infinitesimal deformations
Here, $c_{1}$ is induced by $\Sigma_{\theta}\simeq\Sigma_{t\theta}$, and $c_{2}$ is the morphism $\iota^{{\tt h}}$ in (3.4.1) for $(E,t\theta)$. Let $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. Let us study the harmonic $1$-form ${\tt H}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that ${\tt H}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$. Let $\|\nu\|$ denote the $L^{2}$-norm of $\nu$ on $\Sigma_{\theta}$. Let $g_{X}$ denote a conformal metric of $X$.
@@ 4.3.1 Asymptotically harmonic representative on $X^{\circ}$
$\left|\sqrt{-1}\int_{X\setminus N(D)}h_{t}(F_{\nu},F_{\nu})-\sqrt{-1}\int_{\Sigma_{\theta|X\setminus N(D)}}\nu\wedge\overline{\nu}\right|\\ \leq B_{1}\exp(-B_{2}t)\|\nu\|^{2}.$ (44)
@@
$\sqrt{-1}\int_{\Sigma_{\theta|X\setminus N(D)}}\nu\wedge\overline{\nu}=\sqrt{-1}\int_{X\setminus N(D)}\mathop{\rm tr}\nolimits\bigl(F_{\nu}\cdot F_{\overline{\nu}}\bigr).$
@@
$\sqrt{-1}\int_{X\setminus N(D)}h_{t}(F_{\nu},F_{\nu})=\sqrt{-1}\int_{X\setminus N(D)}\mathop{\rm tr}\nolimits\bigl(F_{\nu}\cdot(F_{\nu})^{\dagger}_{h_{t}}\bigr).$
@@
$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{g_{X}}((\mathop{\rm tr}\nolimits F_{\nu})\mathop{\rm id}\nolimits_{E})=0$.